# A causal inference framework for spatial confounding

### v11

@@ 8 Discussion
There remains much room for further research. For example, unmeasured confounders will often be associated with spatial location even if they are not measurable functions of spatial location. In this scenario, it would be beneficial to quantify the extent to which confounding bias can be reduced by including spatial location in the causal model and/or provide a framework for sensitivity analysis for causal estimates. The assessment of exposures that are perfectly smooth in space (for example, distance to some location such as a factory) is beyond the reach of our current methods due to violation of positivity but may be possible under certain assumptions about relevant scales of variation, as in Paciorek (2010) and Keller and Szpiro (2020). In addition, the mitigation of spatial confounding for areal data has been mostly ignored in this work as some fundamental concepts are distinct in the areal domain; with discrete locations, it is not possible for a confounding variable to be a smooth function, and it is not obvious how to consider an asymptotic regime as the number of areal units is typically fixed.
Supplementary Materials
@@
Assumption S2 ensures that the nuisance functionals are Donsker and is not required if sample splitting is used.
@@
$r\sim\mbox{Unif}(-1,1),X_{i},U_{i}|r,\,\sim_{i.i.d.}N(\*0,\begin{bmatrix}1&r\\ r&1\\ \end{bmatrix}),\,Y_{i}|X_{i},U_{i},r\sim_{i.i.d.}N(X_{i}+U_{i},1)$
@@
OLS / RSR (Hodges and Reich, 2010) (unadjusted estimator) Bootstrap.

### v12

@@ 8 Discussion
There remains much room for further research. For example, unmeasured confounders will often be associated with spatial location even if they are not measurable functions of spatial location. In this scenario, it would be beneficial to quantify the extent to which confounding bias can be reduced by including spatial location in the causal model and/or to provide a framework for sensitivity analysis for causal estimates. The assessment of exposures that are perfectly smooth in space (for example, distance to some location such as a factory) is beyond the reach of our current methods due to violation of positivity but may be possible under certain assumptions about relevant scales of variation, as in Paciorek (2010) and Keller and Szpiro (2020). In addition, the mitigation of spatial confounding for areal data has been mostly ignored in this work as some fundamental concepts are distinct in the areal domain; with discrete locations, it is not possible for a confounding variable to be a smooth function, and it is not obvious how to consider an asymptotic regime as the number of areal units is typically fixed.
Besides spatial confounding, preferential sampling of locations can be another issue which leads to distorted causal effect estimates. Diggle et al. (2010) showed that preferential sampling can bias estimates of spatial structure (variograms) and predictions, but did not study its impact on exposure effect estimates. Pati et al. (2011) considered exposures as well and showed that by modeling the intensity process for sampling of the locations, one can correct for the bias and identify the exposure effect estimate. In Section 3, we present a sufficient set of assumptions under which non-parametric identification (and thereby estimation) of the exposure effect is feasible by using the locations as a proxy for the unmeasured confounder. The framework subsumes certain types of preferential sampling, e.g., ones where the intensity process for sampling depends only on the function $g$ used to define the unmeasured confounder (Assumption 3.3), but not on $X$ or $Y$. However, if the sampling intensity directly depends on $X$ or $Y$, then Assumption (3.5) is violated, and causal identification is no longer feasible. Thus, preferential sampling can sometimes lead to a violation of the assumptions needed for causal identification under spatial confounding, and different adjustment is required. The approaches to adjust for preferential sampling proposed by Diggle et al. (2010); Pati et al. (2011) rely on correct specification of their parametric models, both for the outcome and for the sampling of the locations. Generally, spatial confounding can occur with or without preferential sampling, and vice versa. In the future, more non-parametric framing of preferential sampling and its interplay with unmeasured spatial confounding needs to be studied.
9 Acknowledgments
This work is partially supported by National Institute of Environmental Health Sciences (NIEHS) grant R01ES033739 and Office of Naval Research (ONR) grant N000142112820. The authors would like to thank Heather McBrien for assistance with statistical programming. The authors are also grateful for the use of the facilities at the Joint High Performance Computing Exchange (JHPCE) in the Department of Biostatistics, Johns Hopkins Bloomberg School of Public Health that have contributed to the research results reported within this paper.
Supplementary materials for “A causal inference framework for spatial confounding”
@@ 9 Acknowledgments
Assumption S2.2 ensures that the nuisance functionals are Donsker and is not required if sample splitting is used.
@@
$r\sim\mbox{Unif}(-1,1),X_{i},U_{i}|r,\,\sim_{i.i.d.}N(\mathbf{0},\begin{bmatrix}1&r\\ r&1\\ \end{bmatrix}),\,Y_{i}|X_{i},U_{i},r\sim_{i.i.d.}N(X_{i}+U_{i},1)$
@@
OLS / RSR (Hodges and Reich, 2010) (unadajusted estimator) Bootstrap.